OpenStudy (anonymous):

At the Crab Shack, a party of four ordered crab cakes and another party of four ordered shrimp. Their total bill was $31. However, if only two people ordered the crab cakes and one person ordered the shrimp, the bill would have been $12.25. How much was each order of crab cakes and shrimp? Crab cakes - $4.50;Shrimp - $3.25 Not enough information Crab cakes - $2.04;Shrimp - $4.08 Crab cakes - $3.25;Shrimp - $4.50

OpenStudy (anonymous):

Let c and s be the cost of each crab cake and shrimp dish respectively. Solve the following equations for c and s,\[\{4c+4s=3100,2c+s=1225\}\]where the costs are in pennies. c=450 and s=325 or c=$4.50 and s = $3.25

OpenStudy (anonymous):

so just plug the prices in the equation

OpenStudy (anonymous):

If you want to verify the results, yes. Have you been thought how to solve simultaneous equations?

OpenStudy (anonymous):

xD ??????????? what is that

OpenStudy (anonymous):

Thx but i dont know what simultaneous equations are

OpenStudy (anonymous):

http://www.allaboutcircuits.com/vol_5/chpt_4/11.html

OpenStudy (anonymous):

thx i have more questions like this and your website should help thx again

OpenStudy (anonymous):

Take the second equation and solve for s. By inspection, s = 1225-2c Replace s in the first equation with (1225-2c) 4 c + 4 (1225 - 2 c) = 3100 Can you solve the above equation for c?

OpenStudy (anonymous):

wait what is this i thought we were done with this question

OpenStudy (anonymous):

Well I was going to show how to solve this problem. Thank you for the medal.